OpenStudy (anonymous):

log (x+3)=log(7-4x) how to solve that?

OpenStudy (anonymous):

do you have an iphone

OpenStudy (anonymous):

There is a rule that if you have the same base then the values inside are equal. For example if log(a)=log(b) then it follows that a=b. From their you just solve for x.

OpenStudy (anonymous):

so it will be x+3=7-4x x+4x=7-3 5x=4 x=0.8

OpenStudy (anonymous):

Yeah :)